OpenStudy (i_love_my_nieces):

2, 1, 1/2, 1/4, . . . Arithmetic Geometric Both Neither

OpenStudy (texaschic101):

arithmetic and geometric arithmetic : difference is : add - 1/2 to each term to find the next term geometric : difference is : multiply each term by 1/2 to find the next term

OpenStudy (i_love_my_nieces):

It cannot be both @texaschic101

OpenStudy (texaschic101):

It is both

OpenStudy (i_love_my_nieces):

But it can't

OpenStudy (i_love_my_nieces):

@texaschic101

OpenStudy (mathstudent55):

2, 1, 1/2, 14 Check for arithmetic sequence. 1) Look at the first and second terms: 2, 1 Subtract the first term from the second term: 1 - 2 = -1 2) Look at the second and third terms: 1, 1/2 Subtract the second term from the third term: 1/2 - 1 = -1/2 The two differences are different, so it cannot be an arithmetic sequence.

OpenStudy (mathstudent55):

Now to check for a geometric sequence, follow the same steps above, but divide the terms instead of subtracting them. If you get the same result every time you divide consecutive terms, it is a geometric sequence.

OpenStudy (texaschic101):

your right...I am sorry....not arithmetic

OpenStudy (i_love_my_nieces):

That's what I thought. :)

OpenStudy (texaschic101):

again...I am so sorry :(

OpenStudy (mathstudent55):

2, 1, 1/2, 1/4 Check for geometric sequence: (1)/(2) = 1/2 (1/2)/(1) = 1/2 (1/4)/(1/2) = 1/2 All divisions give you the same answer, so it's geometric.

OpenStudy (i_love_my_nieces):

@mathstudent55 can you help me with a few more?

OpenStudy (mathstudent55):

Yes, but please start a new post for each new question.

OpenStudy (i_love_my_nieces):

Okay, thank you